### v3

# Second Generation 3-DAP 3-Dimensional Assignment Problem

2G
Second Generation
3-DAP
3-Dimensional Assignment Problem
3G
[…]
3GPP
[…]
4G
[…]
5G
[…]
AA
Antenna Array
AC
Admission Control
AD
Attack-Decay
ADC
analog-to-digital converter
ADMM
alternating direction method of multipliers
ADSL
Asymmetric Digital Subscriber Line
AHW
Alternate Hop-and-Wait
AI
Artificial Intelligence
AirComp
over-the-air computation
AirFL
over-the-air federated learning
AMC
Adaptive Modulation and Coding
ANN
artificial neural network
AP
Access Point
APA
Adaptive Power Allocation
ARMA
Autoregressive Moving Average
ARQ
Automatic Repeat Request
ATES
Adaptive Throughput-based Efficiency-Satisfaction Trade-Off
AWGN
additive white Gaussian noise
BAA
Broadband Analog Aggregation
$B$
Branch and Bound
BCD
block coordinate descent
BD
Block Diagonalization
BER
Bit Error Rate
BF
Best Fit
BFD
bidirectional full duplex
BLER
BLock Error Rate
BPC
Binary Power Control
BPSK
Binary Phase-Shift Keying
BRA
Balanced Random Allocation
BS
base station
BSUM
block successive upper-bound minimization
CAP
Combinatorial Allocation Problem
CAPEX
Capital Expenditure
CBF
Coordinated Beamforming
CBR
Constant Bit Rate
CBS
Class Based Scheduling
$C$
Congestion Control
CDF
Cumulative Distribution Function
CDMA
Code-Division Multiple Access
CE
Channel Estimation
CFO
carrier frequency offset
CL
Closed Loop
CLPC
Closed Loop Power Control
CML
centralized machine learning
CNR
Channel-to-Noise Ratio
CNN
Convolutional Neural Network
CP
computation point
CPA
Cellular Protection Algorithm
CPICH
Common Pilot Channel
CoCoA
Communication efficient distributed dual Coordinate Ascent
CoMAC
Computation over Multiple-Access Channels
CoMP
Coordinated Multi-Point
CQI
Channel Quality Indicator
CRM
Constrained Rate Maximization
CRN
Cognitive Radio Network
CS
Coordinated Scheduling
CSI
Channel State Information
CSMA
Carrier Sense Multiple Access
CUE
Cellular User Equipment
D2D
device-to-device
DAC
digital-to-analog converter
DC
direct current
DCA
Dynamic Channel Allocation
DE
Differential Evolution
DFT
Discrete Fourier Transform
DIST
Distance
DL
downlink
DMA
Double Moving Average
DML
Distributed Machine Learning
DMRS
demodulation reference signal
D2DM
D2D Mode
DMS
D2D Mode Selection
DPC
Dirty Paper Coding
DRA
Dynamic Resource Assignment
DSA
Dynamic Spectrum Access
DSGD
Distributed Stochastic Gradient Descent
DSM
Delay-based Satisfaction Maximization
ECC
Electronic Communications Committee
EFLC
Error Feedback Based Load Control
EI
Efficiency Indicator
eNB
Evolved Node B
EPA
Equal Power Allocation
EPC
Evolved Packet Core
EPS
Evolved Packet System
E-UTRAN
Evolved Universal Terrestrial Radio Access Network
ES
Exhaustive Search
FC
Fusion Center
FD
Federated Distillation
FDD
frequency divisionov duplex
FDM
Frequency Division Multiplexing
FDMA
Frequency Division Multiple Access
FedAvg
Federated Averaging
FER
Frame Erasure Rate
$F$
Fast Fading
FL
federated learning
FOMs
first-order methods
FPP-SCA
Feasible Point Pursuit-Successive Convex Approximation
FSB
Fixed Switched Beamforming
FST
Fixed SNR Target
FTP
File Transfer Protocol
GA
Genetic Algorithm
GBR
Guaranteed Bit Rate
GLR
Gain to Leakage Ratio
GOS
Generated Orthogonal Sequence
GPL
GNU General Public License
GRP
Grouping
HARQ
Hybrid Automatic Repeat Request
HD
half-duplex
HMS
Harmonic Mode Selection
HOL
Head Of Line
HSDPA
High-Speed Downlink Packet Access
HSPA
High Speed Packet Access
HTTP
HyperText Transfer Protocol
ICMP
Internet Control Message Protocol
ICI
Intercell Interference
ID
Identification
IETF
Internet Engineering Task Force
ILP
Integer Linear Program
JRAPAP
Joint RB Assignment and Power Allocation Problem
UID
Unique Identification
IID
Independent and Identically Distributed
IIR
Infinite Impulse Response
ILP
Integer Linear Problem
IMT
International Mobile Telecommunications
INV
Inverted Norm-based Grouping
IoT
Internet of Things
IP
Integer Programming
IPMs
interior-point methods
IPv6
Internet Protocol Version 6
ISD
Inter-Site Distance
ISI
Inter Symbol Interference
ITU
International Telecommunication Union
JAFM
joint assignment and fairness maximization
JAFMA
joint assignment and fairness maximization algorithm
JOAS
Joint Opportunistic Assignment and Scheduling
JOS
Joint Opportunistic Scheduling
JP
Joint Processing
JS
Jump-Stay
KKT
Karush-Kuhn-Tucker
L3
Layer-3
LAC
Link Admission Control
LA
Link Adaptation
LC
Load Control
LDC
Learning-Driven Communication
LOS
line of sight
LP
Linear Programming
LTE
Long Term Evolution
LTE-A
Long Term Evolution (LTE)-Advanced
LTE-Advanced
Long Term Evolution Advanced
LRA
Low Rank Approximation
M2M
Machine-to-Machine
MAC
multiple access channel
MAP
maximum a posterior
MANET
Mobile Ad hoc Network
MC
Modular Clock
MCS
Modulation and Coding Scheme
MDB
Measured Delay Based
MDI
Minimum D2D Interference
MF
Matched Filter
MG
Maximum Gain
MH
Multi-Hop
MIMO
Multiple Input Multiple Output
MINLP
mixed integer nonlinear programming
MIP
mixed integer programming
MISO
multiple input single output
ML
machine learning
MLE
maximum likelihood estimator
MLWDF
Modified Largest Weighted Delay First
MME
Mobility Management Entity
MMSE
minimum mean squared error
MOS
Mean Opinion Score
MPF
Multicarrier Proportional Fair
MRA
Maximum Rate Allocation
MR
Maximum Rate
MRC
Maximum Ratio Combining
MRT
Maximum Ratio Transmission
MRUS
Maximum Rate with User Satisfaction
MS
Mode Selection
MSE
Mean Squared Error
MSI
Multi-Stream Interference
MTC
Machine-Type Communication
MTSI
Multimedia Telephony Services over IMS
MTSM
Modified Throughput-based Satisfaction Maximization
MU-MIMO
Multi-User Multiple Input Multiple Output
MU
Multi-User
NAS
Non-Access Stratum
NB
Node B
NCL
Neighbor Cell List
NLP
Nonlinear Programming
NLOS
non-line of sight
NMSE
normalized mean square error
NOMA
Non-Orthogonal Multiple Access
NORM
Normalized Projection-based Grouping
NP
non-polynomial time
NRT
Non-Real Time
NSPS
National Security and Public Safety Services
O2I
Outdoor to Indoor
OFDMA
Orthogonal Frequency Division Multiple Access
OFDM
Orthogonal Frequency Division Multiplexing
OFPC
Open Loop with Fractional Path Loss Compensation
O2I
Outdoor-to-Indoor
OL
Open Loop
OLPC
Open-Loop Power Control
OL-PC
Open-Loop Power Control
OPEX
Operational Expenditure
ORB
Orthogonal Random Beamforming
JO-PF
Joint Opportunistic Proportional Fair
OSI
Open Systems Interconnection
PAIR
D2D Pair Gain-based Grouping
PAPR
Peak-to-Average Power Ratio
P2P
Peer-to-Peer
PC
Power Control
PCI
Physical Cell ID
PDCCH
physical downlink control channel
PDD
penalty dual decomposition
PDF
Probability Density Function
PER
Packet Error Rate
PF
Proportional Fair
P-GW
Packet Data Network Gateway
PL
Pathloss
RLT
reformulation linearization technique
PRB
Physical Resource Block
PROJ
Projection-based Grouping
ProSe
Proximity Services
PS
Parameter Server
PSO
Particle Swarm Optimization
PUCCH
physical uplink control channel
PZF
Projected Zero-Forcing
QAM
quadrature amplitude modulation
QoS
quality of service
QPSK
quadrature phase shift keying
QCQP
quadratically constrained quadratic programming
RAISES
Reallocation-based Assignment for Improved Spectral Efficiency and Satisfaction
RAN
Radio Access Network
RA
Resource Allocation
RAT
Radio Access Technology
RATE
Rate-based
RB
resource block
RBG
Resource Block Group
REF
Reference Grouping
ReLU
rectified linear unit
ReMAC
repetition for multiple access computing
RF
radio frequency
RLC
Radio Link Control
RM
Rate Maximization
RNC
Radio Network Controller
RND
Random Grouping
RRA
Radio Resource Allocation
RRM
Radio Resource Management
RSCP
Received Signal Code Power
RSRP
reference signal receive power
RSRQ
Reference Signal Receive Quality
$R$
Round Robin
RRC
Radio Resource Control
RSSI
received signal strength indicator
RT
Real Time
RU
Resource Unit
RUNE
RUdimentary Network Emulator
RV
Random Variable
SAA
Small Argument Approximation
SAC
Session Admission Control
SCA
successive convex approximation
SCM
Spatial Channel Model
SC-FDMA
Single Carrier - Frequency Division Multiple Access
SD
subgradient descent
S-D
Source-Destination
SDPC
Soft Dropping Power Control
SDMA
Space-Division Multiple Access
SDR
semidefinite relaxation
SDP
semidefinite programming
SeMAC
sequential modulation for AirComp
SeMAC-PA
SeMAC with power adaptation
SER
Symbol Error Rate
SES
Simple Exponential Smoothing
S-GW
Serving Gateway
SGD
Stochastic Gradient Descent
SINR
signal-to-interference-plus-noise ratio
SI
self-interference
SIP
Session Initiation Protocol
SISO
Single Input Single Output
SIMO
Single Input Multiple Output
SIR
Signal to Interference Ratio
SLNR
Signal-to-Leakage-plus-Noise Ratio
SMA
Simple Moving Average
SNR
Signal-to-Noise Ratio
SORA
Satisfaction Oriented Resource Allocation
SORA-NRT
Satisfaction-Oriented Resource Allocation for Non-Real Time Services
SORA-RT
Satisfaction-Oriented Resource Allocation for Real Time Services
SPF
Single-Carrier Proportional Fair
SRA
Sequential Removal Algorithm
SRS
sounding reference signal
SSD
stochastic subgradient descent
SU-MIMO
Single-User Multiple Input Multiple Output
SU
Single-User
SVD
singular value decomposition
SVM
Support Vector Machine
TCP
Transmission Control Protocol
TDD
time division duplex
TDMA
Time Division Multiple Access
TNFD
three node full duplex
TETRA
Terrestrial Trunked Radio
TP
Transmit Power
TPC
Transmit Power Control
TTI
transmission time interval
TTR
Time-To-Rendezvous
TSM
Throughput-based Satisfaction Maximization
TU
Typical Urban
UE
User Equipment
UEPS
Urgency and Efficiency-based Packet Scheduling
UL
uplink
UMTS
Universal Mobile Telecommunications System
URI
Uniform Resource Identifier
URM
Unconstrained Rate Maximization
VR
Virtual Resource
VoIP
Voice over IP
WAN
Wireless Access Network
WCDMA
Wideband Code Division Multiple Access
WF
Water-filling
WiMAX
Worldwide Interoperability for Microwave Access
WINNER
Wireless World Initiative New Radio
WLAN
Wireless Local Area Network
WMMSE
weighted minimum mean square error
WMPF
Weighted Multicarrier Proportional Fair
WPF
Weighted Proportional Fair
WSN
Wireless Sensor Network
WWW
World Wide Web
XIXO
(Single or Multiple) Input (Single or Multiple) Output
ZF
Zero-Forcing
ZMCSCG
Zero Mean Circularly Symmetric Complex Gaussian
Multi-Symbol Digital AirComp via
Modulation Design and Power Adaptation
Xiaojing Yan, , Saeed Razavikia, , Carlo Fischione,
All the authors are with the School of Electrical Engineering and Computer Science KTH Royal Institute of Technology, Stockholm, Sweden (e-mail: {xiay,sraz,carlofi}@kth.se). C. Fischione is also with Digital Futures of KTH.
Abstract
Recently, over-the-air computation (AirComp) leverages the superposition property of wireless channels to enable efficient function computation over a multiple access channel (MAC). However, existing digital AirComp methods either rely on single-symbol modulation, which limits flexibility and robustness, or on multi-symbol extensions that suffer from high complexity or approximation errors. To overcome these limitations, we propose a new multi-symbol modulation framework, termed sequential modulation for AirComp (SeMAC), which encodes each input into a sequence of symbols with distinct constellation diagrams across multiple time slots. This approach increases design flexibility and robustness against channel noise. Specifically, the modulation design is formulated as a non-convex optimization problem and efficiently solved through a successive convex approximation (SCA) combined with stochastic subgradient descent (SSD). For fixed modulation formats, we further develop SeMAC with power adaptation (SeMAC-PA) to adjusts transmit power and phase while preserving the modulation structure. Notably, numerical results show that SeMAC improves computation accuracy by up to 14 dB compared to the existing methods for computing nonlinear functions such as the product function.
I Introduction
With the widespread deployment of Internet of Things (IoT), cellular networks are shifting from “connected everything” in 5G to “connected intelligence” in 6G systems [1]. This shift is driven by the growing need to process distributed data at the network edge devices [2]. In this context, AirComp has emerged as a promising technique that leverages the signal superposition property to compute functions directly during transmission [3]. In particular, AirComp supports 6G intelligence applications such as collaborative sensing and over-the-air federated learning (AirFL), where local model updates are efficiently aggregated across devices to enable global model training [4].
Analog AirComp relies on uncoded transmission, making it vulnerable to noise and fading, and incompatible with modern digital infrastructures. To address these limitations, several digital AirComp methods have been developed, such as one-bit broadband digital aggregation (OBDA), majority-vote frequency-shift keying (FSK), and their asynchronous and non-coherent extensions. However, these methods are often limited to fewer number of transmitters or specific functions [3]. To provide a more general solution, ChannelComp [5] was introduced to enable arbitrary function computation over a MAC by designing digital modulation schemes via optimization, and it adopts a single-symbol modulation design where each input is mapped to a single modulated symbol.
While ChannelComp is compatible with existing digital communication systems, its single-symbol design limits computation accuracy under noisy conditions and its semidefinite relaxation (SDR) solution involves high optimization complexity. As a result, recent works have focused on multi-symbol modulation designs, which provide additional redundancy and improve robustness [6, 7, 8]. For instance, repetition for multiple access computing (ReMAC[6] enhances computation robustness by selectively repeating modulated symbols across multiple time slots, but this repetition coding design further increases computational complexity. Bit-Slicing [8] offers a lower computational complexity design by partitioning bit sequences into slices and mapping each sliced integer to square quadrature amplitude modulation (QAM) symbols for transmission, yet it is mainly effective for the sum function and suffers from approximation errors when extended to nonlinear computations.
Motivated by these limitations, we propose SeMAC, a new multi-symbol modulation framework that encodes each input into a sequence of symbols using distinct constellation diagrams. As opposed to ReMAC which uses an identical constellation pattern across all time slots, SeMAC leverages constellation diversity to achieve higher computation accuracy. The resulting modulation design problem is formulated as a non-convex quadratically constrained quadratic programming (QCQP), which is NP-hard and computationally demanding [9]. To address this complexity challenge, we develop an iterative algorithm based on SCA combined with SSD, which achieves lower complexity than the standard SDR method. In addition, for scenarios where the modulation format is fixed, we propose SeMAC-PA, which adapts the transmit amplitude and phase of the existing modulation patterns, thereby extending the applicability of our approach to hardware-constrained systems. Numerical results show that SeMAC enables robust computation over a noisy MAC and reduces the computation error by up to 14 dB compared to ReMAC, particularly for the product function.
[…]
Figure 1: The overall process of proposed SeMAC via multi-symbol modulation design.
II Preliminaries
II-A Signal Model
Consider a network with $K$ single-antenna nodes transmitting data to a computation point (CP) over a shared channel. Each node $k$ sends a data value $x_{k}\in\mathbb{F}_{Q}$, where $\mathbb{F}_{Q}$ denotes a finite field with $Q=2^{b}$ elements, where $x_{k}$ is represented using $b$ bits. The goal of the CP is to compute any finite-valued function $f(x_{1},\ldots,x_{K})$ over the MACs. Each node encodes its data value $x_{k}$ into a modulated signal $\vec{x}_{k}$ through an encoder $\mathcal{E}_{k}$, such that $\vec{x}_{k}=\mathcal{E}_{k}(x_{k})\in\mathbb{C}$. We assume perfect synchronization of carrier frequency and symbol timing among all the nodes and the CP [^1] In practical systems, techniques such as frame timing and carrier frequency offset estimation can be applied for achieving synchronization [10]., and then the CP receives the superimposed signal as
$\vec{y}=\sum\nolimits_{k=1}^{K}h_{k}p_{k}\vec{x}_{k}+\vec{z},$ (1)
where $h_{k}$ and $p_{k}$ denote the channel coefficient and transmit power for node $k$, and $\vec{z}\sim\mathcal{N}(0,\sigma_{z}^{2})$ represents the additive white Gaussian noise.
Assuming the availability of perfect Channel State Information (CSI), each node applies optimal power control by inverting the channel [11], given by $p_{k}=h_{k}^{*}/|h_{k}|^{2}$. In this regard, we can compensate for channel distortion and ensure coherent signal alignment at the CP. With channel inversion applied, and ignoring the effect of the noise, the received signal $\vec{r}:=\sum\nolimits_{k=1}^{K}\vec{x}_{k}$ yields a finite constellation diagram. Then, the CP applies a tabular function $\mathcal{T}(\cdot)$ to map each constellation point to its corresponding function output.
II-B Overlap Avoidance
Consider a noiseless MAC, and let $f^{(i)}$ and $f^{(j)}$ be two distinct function values corresponding to input sets $x_{1}^{(i)},\dots,x_{K}^{(i)}$ and $x_{1}^{(j)},\dots,x_{K}^{(j)}$, respectively. The aggregated constellation points are given by $\vec{v}^{(i)}:=\sum\nolimits_{k=1}^{K}\vec{x}_{k}^{(i)}$ and $\vec{v}^{(j)}:=\sum\nolimits_{k=1}^{K}\vec{x}_{k}^{(j)}$. To ensure valid computation, destructive overlap between $\vec{v}^{(i)}$ and $\vec{v}^{(j)}$ must be avoided, so that the tabular function $\mathcal{T}(\cdot)$ can uniquely map the constellation point $\vec{v}^{(i)}$ to its corresponding function value $f^{(i)}$. This leads to the following overlap avoidance constraint by a smooth condition [5]:
$|\vec{v}^{(i)}-\vec{v}^{(j)}|\geq\epsilon|f^{(i)}-f^{(j)}|^{2},\quad\forall(i,j)\in\mathcal{M},$ (2)
where $\mathcal{M}:=\{(i,j)\in[M]^{2}|i<j\}$ and $\epsilon>0$ is a positive constant. Specifically, $M$ is the cardinality of the function range, and $\tilde{M}$ is the number of index pair in set $\mathcal{M}$.
III System Model
III-A Multi-Symbol Transmission
Extending the signal model in II-A, SeMAC enables multi-symbol transmission by encoding each input value $x_{k}$ at node $k$ into a sequence of $L$ modulated symbols using a modulation encoder, expressed as $\mathcal{E}_{k}(x_{k})=[\vec{x}_{k,1},\dots,\vec{x}_{k,L}]\in\mathbb{C}^{1\times L}$. These symbols are then transmitted sequentially over $L$ consecutive time slots, with $\vec{x}_{k,\ell}$ corresponding to time slot $\ell$. Consequently, the superimposed signal received at the CP in each time slot is given by:
$\vec{y}_{\ell}=\sum\nolimits_{k=1}^{K}h_{k,\ell}p_{k,\ell}\vec{x}_{k,\ell}+\vec{\tilde{z}}_{\ell},\quad\forall~\ell\in[L],$ (3)
where $h_{k,\ell}$ and $p_{k,\ell}$ denote the channel coefficient and transmit power of node $k$ at time slot $\ell$, and $\vec{\tilde{z}}_{\ell}\sim\mathcal{N}(0,\sigma_{z}^{2})$ is the additive white Gaussian noise at time slot $\ell$. Similarly, following the optimal power control policy [11], the received signal can be written as:
$\vec{y}_{\ell}=\sum\nolimits_{k=1}^{K}\vec{x}_{k,\ell}+\vec{\tilde{z}}_{\ell},\quad\forall~\ell\in[L].$ (4)
For each node $k$, we define the modulation matrix as $\bm{X}_{k}\in\mathbb{C}^{Q\times L}$, where $\bm{X}_{k}]_{(q,\ell)}:=\vec{x}_{k,\ell}^{(q)}$ denotes the signal transmitted at time slot $\ell$ when the input is $x_{k}^{(q)}\in\mathbb{F}_{Q}$. In this matrix, each row corresponds to an input value and contains its encoded modulation sequence $\bm{x}_{k}^{(q)}:=\mathcal{E}_{k}(x_{k}^{(q)})=[\vec{x}_{k,1}^{(q)},\ldots,\vec{x}_{k,L}^{(q)}]$. Meanwhile, each column $\bm{x}_{k,\ell}:=[\vec{x}_{k,\ell}^{(1)},\ldots,\vec{x}_{k,\ell}^{(Q)}]^{\mathsf{T}}\in\mathbb{C}^{Q\times 1}$ defines the modulation pattern used at time slot $\ell$, containing the $Q$ complex symbols assigned to all possible input values. Without loss of generality, each modulation vector follows a unit norm, i.e., $\|\bm{x}_{k,\ell}\|_{2}^{2}\leq 1$, $\forall k\in[K]$ and $\forall\ell\in[L]$.
Furthermore, we construct the modulation matrix $\bm{X}:=[\bm{X}_{1}^{\mathsf{T}},\dots,\bm{X}_{K}^{\mathsf{T}}]^{\mathsf{T}}\in\mathbb{C}^{N\times L}$, where $N:=QK$, by vertically stacking the modulation matrices of all $K$ nodes. Give the input combination $x_{1}^{(i)},\dots,x_{K}^{(i)}$, we define a binary vector $\bm{a}_{i}\in\{0,1\}^{N}$ which consists of $K$ one-hot blocks of length $Q$, each selecting the encoded modulation symbol at one node. Therefore, the resultant modulation sequence associated with the function output $f^{(i)}$ is given by $\bm{v}^{(i)}=\bm{a}_{i}\bm{X}\in\mathbb{C}^{L}$.
III-B Multi-Symbol Estimation and Decoding
After receiving the aggregated signal over the MAC, the CP needs to estimate the transmitted constellation sequence, and then maps the estimated sequence to the output of function $f$. More precisely, in each time slot $\ell$, a maximum likelihood estimator partitions the constellation diagram, which consists of all possible constellation points $\{\vec{v}_{\ell}^{(1)},\ldots,\vec{v}_{\ell}^{(M)}\}$ along with their corresponding Voronoi cells $\{\mathcal{V}_{1,\ell},\ldots,\mathcal{V}_{M,\ell}\}$, and maps the received symbol $\vec{y}_{\ell}$ to the closest constellation point as $\vec{v}_{\ell}=\arg\min_{i}\|\vec{y}_{\ell}-\vec{v}^{(i)}_{\ell}\|_{2}^{2}$. Then, the recovered function output $\hat{f}$ is computed via the tabular function $\hat{f}=\sum\nolimits_{j=1}^{M}\mathcal{T}_{j}(\bm{v})$, where $\mathcal{T}_{j}(\cdot)$ is an indicator function:
$\mathcal{T}_{j}(\bm{v}):=\begin{cases}\hat{f}^{(j)},&\quad\text{if }\vec{v}_{\ell}\in\mathcal{V}_{j,\ell},\text{ }\forall\ell\in[L],\\ 0,&\quad\textrm{otherwise}.\end{cases}$ (5)
To ensure error-free computation, the tabular function must uniquely associate each received constellation sequence with its corresponding function value. In contrast to ChannelComp, where decoding relies on individual symbols, correct recovery in SeMAC depends on the received modulation sequence. Thus, any two sequences $\bm{v}_{i}$ and $\bm{v}_{j}$ must be separated in Euclidean space when their function outputs $f^{(i)}$ and $f^{(j)}$ are different. This requirement leads to a generalized overlap avoidance constraint for designing multi-symbol modulation:
$\|(\bm{a}_{i}-\bm{a}_{j})^{\mathsf{T}}\bm{X}\|_{2}^{2}\geq\epsilon|f^{(i)}-f^{(j)}|,\quad\forall(i,j)\in\mathcal{M},$ (6)
where $\|\cdot\|_{2}$ denotes the $\ell_{2}$-norm. Note that if the condition in (6) is not satisfied, the corresponding Voronio regions are merged, and we consider the average of the associated outputs as the final function output [5]. The overall SeMAC process is shown in Figure 1.
IV Modulation Design and Power Adaptation
In this section, we consider a scenario where the network allocates up to $L$ time slots, with $Q>L$. When modulation patterns are configurable, we employ SeMAC to design the modulation. In contrast, when modulation is fixed, we adopt SeMAC-PA, which adapts the transmit power and phase to reshape the received sequence of the constellation.
IV-A Modulation Design With Channel Inversion
In this subsection, the modulation scheme is designed under the assumption of channel inversion power control. To avoid constellation overlaps as in (6), we pose the following feasibility problem to enable robust computation, subject to the norm power constraint:
$\mathcal{P}_{0}:\quad$ ${\rm find}\quad\bm{X}$
${\rm s.t.}\quad$ ${\rm Tr}(\bm{X}^{\mathsf{H}}\bm{D}_{ij}\bm{X})\geq\Delta f_{ij},\quad\forall(i,j)\in\mathcal{M},$ (7a)
$\|\bm{x}_{k,\ell}\|_{2}^{2}\leq 1,\quad\forall k\in[K],\quad\forall\ell\in[L],$ (7b)
where $\Delta f_{ij}:=\epsilon|f^{(i)}-f^{(j)}|$ and $\bm{D}_{ij}:=(\bm{a}_{i}-\bm{a}_{j})(\bm{a}_{i}-\bm{a}_{j})^{\mathsf{T}}$. However, $\mathcal{P}_{0}$ is a non-convex QCQP, and solving it is NP-hard [5]. A common approach to tackle such problems is the SDR [6], which uses matrix lifting coupled with rank relaxation to relax the problem into a convex semidefinite programming (SDP). While effective for small instances, SDR consumes potentially high complexity for large $K$ and $Q$. To overcome this, we consider a scalable alternative based on SCA to approximate the original problem via a sequence of convex subproblems [12]. Specifically, we first formulate the feasibility problem into an unconstrained minimization problem by hinge-loss penalites:
$\mathcal{P}_{1}:\quad\min_{\bm{X}\in\mathcal{X}}\quad\frac{1}{\tilde{M}}\sum\nolimits_{(i,j)\in\mathcal{M}}g_{ij}(\bm{X}),$ (8)
where $g_{ij}(\bm{X}):=\left[\Delta f_{ij}-\operatorname{Tr}(\bm{X}^{\mathsf{H}}\bm{D}_{ij}\bm{X})\right]^{+}$ penalizes constraint violations and $\mathcal{X}:=\{\|\bm{x}_{k,\ell}\|_{2}^{2}\leq 1,\forall k\in[K],\forall\ell\in[L]\}$ is a convex set of power-constrained vectors. Since $g_{ij}(\bm{X})$ is non-convex, we then approximate it by a convex majorizer at a reference point $\tilde{\bm{X}}^{(s)}$, leading to the following function:
$h_{ij}^{(s)}(\bm{X}):=$ $\left[\Delta f_{ij}-2\Re\{\operatorname{Tr}[(\tilde{\bm{X}}^{(s)})^{\mathsf{H}}\bm{D}_{ij}\bm{X}]\}\right$.
$\left.+{\rm Tr}[(\tilde{\bm{X}}^{(s)})^{\mathsf{H}}\bm{D}_{ij}\tilde{\bm{X}}^{(s)}]\right]^{+}.$ (9)
Hence, at SCA iteration $s$, we solve the following convex subproblem $\mathcal{P}_{2}$, whose solution $\bm{\tilde{X}}^{(s+1)}$ serves as the reference point for the next iteration:
$\mathcal{P}_{2}:\bm{\tilde{X}}^{(s+1)}\in\arg\min_{\bm{X}\in\mathcal{X}}\frac{1}{\tilde{M}}\sum\nolimits_{(i,j)\in\mathcal{M}}h_{ij}^{(s)}(\bm{X}).$ (10)
Algorithm 1 SCA-SSD for solving Problem $\mathcal{P}_{0}$
1:Init: SCA iterations $S$, subgradient steps $T$, step sizes $\{\alpha_{t}\}_{t=0}^{T-1}$. Randomly generate an initial point $\bm{\tilde{X}}^{(0)}\in\mathcal{X}$.
2:for $s=0,1,\dots,S$ do
3:for $t=0,1,\dots,T-1$ do
4:  Sample a single pair $(i_{t},j_{t})$ uniformly from $\mathcal{M}$
5:  $\bm{G}^{(t)}=\partial h^{(s)}_{i_{t}j_{t}}\!\left(\bm{X}^{(t)}\right)$
6:  $\bm{Y}^{(t+1)}=\bm{X}^{(t)}-\alpha_{t}\,\bm{G}^{(t)}$
7:  $\bm{X}^{(t+1)}=\Pi_{\mathcal{X}}\!\left(\bm{Y}^{(t+1)}\right)$
8:end for
9: Set $\bm{\tilde{X}}^{(s+1)}=\bm{X}^{(T)}$
10:end for
11:Output: $\hat{\bm{X}}\leftarrow\bm{\tilde{X}}^{(S+1)}$
Instead of solving $\mathcal{P}_{2}$ by interior-point methods, which consumes high computation cost, we can determine an approximate solution by using the SSD approach. Specifically, at each step $t$, we randomly draw the index pair $(i_{t},j_{t})$ from set $\mathcal{M}$ and apply the following update rule:
$\bm{Y}^{(t+1)}$ $=\bm{X}^{(t)}-\alpha_{t}\cdot\partial h_{i_{t}j_{t}}^{(s)}(\bm{X}^{(t)}),$ (11a)
$\bm{X}^{(t+1)}$ $=\Pi_{\mathcal{X}}(\bm{Y}^{(t+1)}).$ (11b)
where $\alpha_{t}$ is the step size and the subgradient $\partial h_{ij}^{(s)}(\bm{X}^{(t)})$ at the point $\bm{X}^{(t)}$ is computed as:
$\partial h_{ij}^{(s)}(\bm{X}^{(t)})=\begin{cases}-2\bm{D}_{ij}\bm{\tilde{X}}^{(s)},&\text{if }h_{ij}(\bm{X}^{(t)})>0\\ \bm{0},&\textrm{otherwise}.\end{cases}$ (12)
Moreover, $\Pi_{\mathcal{X}}(\cdot)$ denotes the projection onto $\mathcal{X}$ given by:
$\bm{x}_{k,\ell}:=\begin{cases}\bm{y}_{k,\ell},&\text{if }\|\bm{y}_{k,\ell}\|_{2}\leq 1,\\ \frac{\bm{y}_{k,\ell}}{\|\bm{y}_{k,\ell}\|_{2}},&\text{otherwise}.\end{cases}$ (13)
The complete process is summarized in Algorithm 1, which integrates the SCA with SSD updates.
Remark 1.
Note that solving the standard SDR problems require $\mathcal{O}(\tilde{M}^{4}N^{0.5}L^{0.5})$ operations due to matrix lifting [6], which becomes computationally demanding for large $K$ or $Q$. In contrast, the proposed SCA–SSD algorithm samples only one constraint per update, leading to per-iteration complexity $\mathcal{O}(NL)$ and total cost $\mathcal{O}(STNL)$. This linear scaling avoids dependence on $\tilde{M}$ and thereby ensures scalability in larger networks. Furthermore, with a diminishing step-size schedule, the SSD updates follow the expected descent direction and the framework converges in expectation to a stationary point [13].
IV-B Power Adaptation of Existing Modulation
Due to hardware constraints, modifying the existing modulation format at each device is often not feasible. Moreover, channel inversion is sensitive to deep fades and may lead to excessive power consumption under realistic fading conditions. To address this, we propose SeMAC-PA, an extension of the E-ChannelComp framework [5], to enable realible function computation with fixed modulation through power and phase adaptation across multiple time slots.
Given the signal model in (3), we define $\bm{p}_{\ell}:=[p_{1,\ell},\ldots,p_{K,\ell}]^{\mathsf{T}}\in\mathbb{C}^{K}$ and $\bm{h}_{\ell}:=[h_{1,\ell},\ldots,h_{K,\ell}]^{\mathsf{T}}\in\mathbb{C}^{K}$ as the transmit power vector and channel vector at time slot $\ell$, respectively. Let $\tilde{\bm{x}}_{\ell}$ denotes the modulation vector containing a fixed modulation pattern, i.e., QAM. In the presence of fading, the received constellation point corresponding to the function output $f^{(i)}$ at time slot $\ell$ is then given by:
$\tilde{\bm{v}}^{(i)}_{\ell}=\bm{a}_{i}^{\mathsf{T}}\left(\text{diag}(\bm{h}_{\ell})\otimes\bm{I}_{Q}\right)\text{diag}(\tilde{\bm{x}}_{\ell})\left(\bm{I}_{K}\otimes\mathds{1}_{Q}\right)\bm{p}_{\ell},$ (14)
where $\otimes$ denotes the Kronecker product. $\bm{I}_{Q}$ and $\bm{I}_{K}$ are identity matrices of size $Q\times Q$ and $K\times K$. $\mathds{1}_{Q}$ is a $Q\times 1$ vector with all elements equal to one, and the operator $\text{diag}(\cdot)$ constructs a diagonal matrix. For simplicity, we let $\bm{B}_{\ell}:=\left(\text{diag}(\bm{h}_{\ell})\otimes\bm{I}_{Q}\right)\text{diag}(\tilde{\bm{x}}_{\ell})(\bm{I}_{K}\otimes\mathds{1}_{Q})$. Accordingly, we formulate the following QCQP to determine the minimum power vectors $\bm{p}_{\ell}$ that satisfiy the constraints in (6):
$\mathcal{P}_{3}:\quad$ $\min_{\{\bm{p}_{\ell}\}\in\mathcal{P}}\quad\sum\nolimits_{\ell=1}^{L}\|\bm{p}_{\ell}\|_{2}^{2}$
${\rm s.t.}\quad$ $\sum\nolimits_{\ell=1}^{L}\bm{p}_{\ell}^{\mathsf{H}}\bm{C}_{\ell}^{ij}\bm{p}_{\ell}\geq\Delta f_{ij},\quad\forall(i,j)\in\mathcal{M},$ (15a)
where $\bm{C}_{\ell}^{ij}:=\bm{B}_{\ell}^{\mathsf{H}}(\bm{a}_{i}-\bm{a}_{j})(\bm{a}_{i}-\bm{a}_{j})^{\mathsf{T}}\bm{B}_{\ell}$ and $\mathcal{P}:=\{\bm{p}_{\ell}\in\mathbb{C}^{K},\ell\in[L]\}$. In a similar manner to $\mathcal{P}_{0}$, we translate the quadratic constraints in $\mathcal{P}_{3}$ into hinge-loss penalties:
$u_{ij}(\{\bm{p}_{\ell}\})=\left[\Delta f_{ij}-\sum\nolimits_{\ell=1}^{L}\bm{p}_{\ell}^{\mathsf{H}}\bm{C}_{\ell}^{ij}\bm{p}_{\ell}\right]^{+},\quad\forall(i,j)\in\mathcal{M},$
which measure the constraint violations. At each SCA iteration $s$, we linearize these non-convex terms at the reference points $\{\bm{\tilde{p}}_{\ell}^{(s)}\}$, leading to the following convex subproblem:
$\mathcal{P}_{4}:\min_{\{\bm{p}_{\ell}\}\in\mathcal{P}}\sum\nolimits_{\ell=1}^{L}\bm{p}_{\ell}^{\mathsf{H}}\bm{I}_{K}\bm{p}_{\ell}+\frac{1}{\tilde{M}}\sum_{(i,j)\in\mathcal{M}}w_{ij}^{(s)}(\{\bm{p}_{\ell}\}),$ (16)
where the convex majorizer is defined as:
$w_{ij}^{(s)}(\{\bm{p}_{\ell}\}):=$ $\left[\Delta f_{ij}-\sum\nolimits_{\ell=1}^{L}2\Re\{(\bm{\tilde{p}}_{\ell}^{(s)})^{\mathsf{H}}\bm{C}_{\ell}^{ij}\bm{p}_{\ell}\}\right$.
$\left.+\sum\nolimits_{\ell=1}^{L}(\bm{\tilde{p}}_{\ell}^{(s)})^{\mathsf{H}}\bm{C}_{\ell}^{ij}\bm{\tilde{p}}_{\ell}^{(s)}\right]^{+}.$ (17)
For the employment of SSD, a subgradient of $w_{ij}^{(s)}(\{\bm{p}_{\ell}\})$ with respect to each variable $\bm{p}_{\ell}$ is given by
$\partial_{\bm{p}_{\ell}}w_{ij}^{(s)}\left(\{\bm{p}_{\ell}\}\right)=\begin{cases}-2\bm{C}_{\ell}^{ij}\bm{\tilde{p}}_{\ell}^{(s)},&\text{if }w_{ij}^{(s)}\left(\{\bm{p}_{\ell}\}\right)>0,\\ \bm{0},&\text{otherwise}.\end{cases}$ (18)
Overall, the complete procedure for solving $\mathcal{P}_{3}$ is summarized in Algorithm 2.
Algorithm 2 SCA-SSD for solving Problem $\mathcal{P}_{3}$
1:Init: SCA iterations $S$, subgradient steps $T$, step sizes $\{\alpha_{t}\}_{t=0}^{T-1}$. Randomly generate initial points $\{\bm{\tilde{p}}_{\ell}^{(0)}\}\in\mathcal{P}$.
2:for $s=0,1,\dots,S$ do
3:for $t=0,1,\dots,T-1$ do
4:  Sample a single pair $(i_{t},j_{t})$ uniformly from $\mathcal{M}$
5:  $\bm{G}_{\ell}^{(t)}=2\bm{p}_{\ell}^{(t)}+\partial_{\bm{p}_{\ell}}w_{ij}^{(s)}(\{\bm{p}_{\ell}^{(t)}\}),\forall\ell\in[L]$
6:  $\bm{p}_{\ell}^{(t+1)}=\bm{p}_{\ell}^{(t)}-\alpha_{t}\,\bm{G}^{(t)}_{\ell},\forall\ell\in[L]$
7:end for
8: Set $\bm{\tilde{p}}_{\ell}^{(s+1)}=\bm{p}_{\ell}^{(T)},\forall\ell\in[L]$
9:end for
10:Output: $\hat{\bm{p}_{\ell}}\leftarrow\bm{\tilde{p}}_{\ell}^{(S+1)},\forall\ell\in[L]$
V Numerical Experiments
In this section, we evaluate the performance of SeMAC as well as its power adaptive scheme SeMAC-PA, and make a comparison with ChannelComp, ReMAC and Bit-Slicing. Specifically, the performance is evaluated using the normalized mean square error (NMSE) metric, defined as:
$\text{\acs{NMSE}}:=\frac{\sum\nolimits_{j=1}^{N_{s}}|f^{(i)}-\hat{f}_{j}^{(i)}|^{2}}{N_{s}|f_{{\rm max}}-f_{{\rm min}}|^{2}},$ (19)
where $N_{s}=100$ is the number of Monte Carlo trials, $f^{(i)}$ is the desired function value, $\hat{f}^{(i)}_{j}$ is the estimated value in the $j$-th Monte Carlo trial. $f_{{\rm max}}$ and $f_{{\rm min}}$ denote the maximum and minimum values of the function output, respectively. The fading coefficients $h_{k,\ell}$ are modeled as independent complex Gaussian random variables, i.e., $h_{k,\ell}\sim\mathcal{CN}(0,1)$. For the SSD updates, we adopt a diminishing step-size schedule $\alpha_{t}=\alpha_{0}/\sqrt{t+1}$, where $\alpha_{0}=0.5$.
V-A Performance of SeMAC
For the first experiment, we consider a network of $K=8$ nodes and evaluate the computation of the sum function $f=\sum_{k=1}^{K}x_{k}$ and the product function $f=\prod_{k=1}^{K}x_{k}$, where $x_{k}\in{1,2,3,4}$. The results in Figure 2 show that the NMSE consistently decreases as the Signal-to-Noise Ratio (SNR) increases. In addition, increasing the number of time slots from $L=1$ to $L=2$ improves the computation accuracy by providing greater flexibility in constellation design. Moreover, our proposed SCA–SSD method generates solutions that achieve lower NMSE than the SDR-based method, as it directly refines solutions via minimizing a hinge-loss penalization within the original problem space, thereby avoiding the relaxation errors introduced by lifting and randomization in SDR.
[…]
Figure 2: Performance of SeMAC in terms of NMSE across SNRs using the SDR baseline and the proposed SCA–SSD algorithm.
[…]
Figure 3: Performance comparison among ReMAC, SeMAC, ChannelComp and Bit-Slicing with $K=4$ nodes for computing the product function.
V-B Comparison to ChannelComp, ReMAC and Bit-Slicing
In this subsection, we compare ChannelComp with repetition, SeMAC, ReMAC, and Bit-Slicing for computing the product function with $K=4$ nodes over $L=2$ time slots, where input values are selected from $x_{k}\in\{1,2,\ldots,16\}$. As shown in Figure 3, by exploiting diverse modulation patterns across time slots, SeMAC achieves lower NMSE by approximately 14 dB compared to ReMAC. While ChannelComp with repetition and ReMAC offer similar performance, ReMAC achieves greater energy efficiency by selectively repeating modulated symbols. Moreover, Bit-Slicing yields the highest NMSE, due to approximation errors introduced during the pre- and post-processing required to support nonlinear functions, as previously analyzed in [6].
V-C Performance of SeMAC-PA
We analyze the performance of SeMAC-PA under Rayleigh fading with $K=8$ and QAM modulation of orders $Q=4$ and $Q=16$ for computing the sum function over $L=2$ time slots. Figure 4 shows that NMSE decreases with increasing SNR, demonstrating the effectiveness of power adaptation. However, SeMAC-PA relies on fixed QAM constellations, making it less effective than the optimized modulation design in SeMAC under channel inversion power control.
[…]
Figure 4: Performance of SeMAC-PA is evaluated across various SNRs in terms of NMSE with $K=8$ nodes.
VI Conclusion
In this letter, we proposed SeMAC, a multi-symbol modulation framework for digital AirComp that enables reliable function computation through distinct modulation patterns across multiple time slots. To support scenarios with fixed modulation, we introduced SeMAC-PA, which adapts transmit power and phase to reshape the received constellation. One possible future work is to extend SeMAC to scenarios with imperfect CSI and investigate its integration with machine learning techniques for adaptive modulation design.